The number density in a container of neon gas is 4.50×1025 m−3 . The atoms are moving with an rms speed of 625 m/s .

What is the pressure inside the container?

What is the temperature inside the container?
Answer #1

Molecular mass of Neon gas = 20.18 u = 20.18 x 1.66 x 10^-27 = 3.35 x 10^-26 kg

A)

Pressure, P = (1/3) (4.5 x 10^25)(3.35 x 10^-26)(625^2) = 1.96 x 10^5 Pa

B)

Temperature, T = (3.35 x 10^-26)(625^2)/3 x 1.38 x 10^-23 = 316.09 K

Comment in case any doubt please rate my answer ....